# Matroid Products via Submodular Coupling

### v1

@@ 4.1 Preparations
Let $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$ be two set algebras. The functions $\varphi_{1}\colon\mathcal{B}_{1}\to\mathbb{R}$ and $\varphi_{2}\colon\mathcal{B}_{2}\to\mathbb{R}$ are isomorphic if there exists a bijection $f\colon J_{1}\to J_{2}$ such that $f$ and $f^{-1}$ are measurable and $\varphi_{2}(f(X))=\varphi_{1}(X)$ for all $X\in\mathcal{B}_{1}$. By a property of set functions, we mean an isomorphism invariant class of set functions. We call a property $P$ of set functions finitary if a function $\varphi\colon\mathcal{B}\to\mathbb{R}$ is in $P$ if and only if $\varphi/\mathcal{Q}$ is in $P$ for any finite measurable partition $\mathcal{Q}$ of $J$. A finitary property $P$ is closed if for any measurable space $(J,\mathcal{B})$ and any finite measurable partition of $J$, the set of functions $\{f\colon\mathcal{B}\to\mathbb{R}\mid f/\mathcal{Q}\in P\}$ is closed. We call a finitary property $P$ extendable if for any $\varphi\colon 2^{[n]}\to\mathbb{R}$ with property $P$ and any finite, pairwise disjoint sets $A_{1},A_{2},\ldots,A_{n}$ with union $A$, there exists a function $\psi\colon 2^{A}\to\mathbb{R}$ with property $P$ that extends $\varphi$ in the sense that that $\varphi(S)=\psi(\bigcup_{i\in S}A_{i})$ for all $S\subseteq[n]$. We call a finitary property $P$ liftable if any two finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. Note that liftability implies that any finite number of finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. We say that property $P$ is $b$-bounded for some $b\in\mathbb{R}_{+}$ if the range of any function with property $P$ is contained in $[-b,b]$. Finally, a normalized function with property $P$ is called universal for property $P$, if it contains all finite, normalized functions with property $P$ as a quotient.
We need the following well-known lemma (which follows from e.g. [24, Problem 4.43]) and include a proof for completeness.
@@ Acknowledgment.
The authors are grateful to Miklós Abért, Márton Borbényi, Máté András Pálfy and László Márton Tóth for helpful discussions. András Imolay was supported by the Rényi Doctoral Fellowship of the Rényi Institute and by the EKÖP-24 University Excellence Scholarship Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.

### v2

@@ 4.1 Preparations
Let $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$ be two set algebras. The functions $\varphi_{1}\colon\mathcal{B}_{1}\to\mathbb{R}$ and $\varphi_{2}\colon\mathcal{B}_{2}\to\mathbb{R}$ are isomorphic if there exists a bijection $f\colon J_{1}\to J_{2}$ such that $f$ and $f^{-1}$ are measurable and $\varphi_{2}(f(X))=\varphi_{1}(X)$ for all $X\in\mathcal{B}_{1}$. By a property of set functions, we mean an isomorphism invariant class of set functions. We call a property $P$ of set functions finitary if a function $\varphi\colon\mathcal{B}\to\mathbb{R}$ is in $P$ if and only if $\varphi/\mathcal{Q}$ is in $P$ for any finite measurable partition $\mathcal{Q}$ of $J$. A finitary property $P$ is closed if for any measurable space $(J,\mathcal{B})$ and any finite measurable partition of $J$, the set of functions $\{f\colon\mathcal{B}\to\mathbb{R}\mid f/\mathcal{Q}\in P\}$ is closed. We call a finitary property $P$ extendable if for any $\varphi\colon 2^{[n]}\to\mathbb{R}$ with property $P$ and any finite, pairwise disjoint sets $A_{1},\ldots,A_{n}$ with union $A$, there exists a function $\psi\colon 2^{A}\to\mathbb{R}$ with property $P$ that extends $\varphi$ in the sense that that $\varphi(S)=\psi(\bigcup_{i\in S}A_{i})$ for all $S\subseteq[n]$. We call a finitary property $P$ liftable if any two finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. Note that liftability implies that any finite number of finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. We say that property $P$ is $b$-bounded for some $b\in\mathbb{R}_{+}$ if the range of any function with property $P$ is contained in $[-b,b]$. Finally, a normalized function with property $P$ is called universal for property $P$, if it contains all finite, normalized functions with property $P$ as a quotient.
We need the following well-known lemma (which follows from e.g. [25, Problem 4.43]) and include a proof for completeness.
@@ Acknowledgment.
The authors are grateful to Miklós Abért, Márton Borbényi, Máté András Pálfy and László Márton Tóth for helpful discussions. András Imolay was supported by the Rényi Doctoral Fellowship of the Rényi Institute and by the EKÖP-24 University Excellence Scholarship Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant numbers ADVANCED 150556 and ELTE TKP 2021-NKTA-62, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115. This work was supported in part by EPSRC grant EP/X030989/1.
Data availability
No data are associated with this article. Data sharing is not applicable to this article.